# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v2

@@ Appendix D Remez algorithm
     $\begin{bmatrix}a_{k}\\ b_{k}\\ c_{k}\\ E_{k}\end{bmatrix}=\begin{bmatrix}\ell&\ell^{3}&\ell^{5}&1\\ x_{1}^{(k)}&(x_{1}^{(k)})^{3}&(x_{1}^{(k)})^{5}&-1\\ x_{2}^{(k)}&(x_{2}^{(k)})^{3}&(x_{2}^{(1)})^{5}&1\\ u&u^{3}&u^{5}&-1\end{bmatrix}^{-1}\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$
     $x_{1}^{(k+1)}=\sqrt{\frac{-3b_{k}-\sqrt{9b_{k}^{2}-20a_{k}c_{k}}}{10c_{k}}},x_{2}^{(k+1)}=\sqrt{\frac{-3b_{k}+\sqrt{9b_{k}^{2}-20a_{k}c_{k}}}{10c_{k}}}$
@@ Appendix E Initialization for Matrices with Large Spectral Gaps
In Section 4, we constructed a sequence of polynomials that is adapted to the range of the singular values $[\ell,u]$. Assuming nothing else about the input, these polynomials are optimal because they provide a good approximation to $1$ across the entire interval. However, in many applications, the spectrum has large gaps; that is, there are several large outlying singular values that are well-separated from the rest. For these matrices, it is not necessary for the polynomial to be accurate on the entire interval $[\ell,u]$, only on the range of the small singular values plus a few other isolated points. In this section, we take advantage of this structure to accelerate our method by preprocessing the matrix to eliminate the largest singular values.
The first step is to find small intervals containing each of these large singular values. To find lower bounds, we use subspace iteration, which is a generalization of the power method that approximates multiple singular values simultaneously. Fix $k$, the number of singular values we wish to eliminate. Letting $\sigma_{1}\geq\cdots\geq\sigma_{n}$ denote the singular values of $\bm{M}$, subspace iteration produces estimates $\tilde{\sigma}_{1}\geq\cdots\geq\tilde{\sigma}_{k}$ satisfying $\sigma_{i}\geq\tilde{\sigma}_{i}$ for all $i\in 1,\ldots,k$ [^4] Let $\bm{Q}_{0}\in\mathbb{R}^{n\times k}$ be a random matrix with orthonormal columns and define $\bm{Q}_{t+1},\bm{R}_{t+1}=\mathtt{qr}\left(\bm{M}^{\top}\bm{M}\bm{Q}_{t}\right)$, where $\mathtt{qr}$ is the QR decomposition. Subspace iteration outputs the singular values $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{k}$ of $\bm{M}\bm{Q}_{T}$, $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{k}$. By the Cauchy interlacing theorem, $\tilde{\sigma}_{k}\leq\sigma_{k}$. To find upper bounds on each $\sigma_{i}$, we can use the fact that $\|\bm{M}\|_{\text{F}}^{2}=\sum_{j=1}^{n}\sigma_{j}^{2}$ as follows:
@@
The second step is to find an odd polynomial that well-approximates the constant function on each of these intervals and on the tail simultaneously. For simplicity, we treat only the $k=1$ case here. Assume that $\bm{M}$ is normalized to $\|\bm{M}\|_{\text{F}}=1$ and let $z=\tilde{\sigma}_{1}$ be the lower bound produced by subspace iteration (which reduces to the power method in this case). Then (19) gives $\sigma_{1}\in[z,1]$ and $\sigma_{2},\ldots,\sigma_{n}\leq\sqrt{1-z^{2}}$. Assume that these intervals do not overlap, that is, $\sqrt{1-z^{2}}\leq z\iff z\geq 1/\sqrt{2}$. Then we construct the unique odd cubic polynomial $p(x)=ax+bx^{3}$ that satisfies $p(\sqrt{1-z^{2}})=1$ and $p(z)=1$ by setting
@@
Because $p(0)=0$ and $p$ has at most one local extremum on $\mathbb{R}_{\geq 0}$, these conditions immediately guarantee that $p$ is concave-increasing on $[0,\sqrt{1-z^{2}}]$, so it must lie above the line $x\mapsto x/\sqrt{1-z^{2}}$. Furthermore, $p$ is decreasing on $[\sigma_{1},1]$, so it maps $\sigma_{1}\in[z,1]$ to $[p(1),1]$. By minimizing $p(1)$ over all valid $z$ (that is, over the interval $z\in[1/\sqrt{2},1]$), one can further show that $p(1)>1/\sqrt{2}$, so $\sigma_{1}$ cannot be decreased very much by applying $p$. Thus, the largest singular value of $p(\bm{M})$ is still at most $1$, while the smaller singular values have increased by a potentially large factor of $1/\sqrt{1-z^{2}}$. When there is a large outlying singular value, $z$ is close to $1$ and this initialization scheme makes much more progress than a standard iteration of PolarExpress would have.
@@ Appendix F Fast Polynomial Iteration for Rectangular Matrices
     $\bm{R}_{t}=\bm{Q}_{t-1}^{\top}\bm{Y}\bm{Q}_{t-1}$triangleright $2n^{3}$
     $\bm{Q}_{t}=\bm{Q}_{t-1}h_{t}(\bm{R}_{t})$triangleright Horner’s rule: $\deg(h_{t})\cdot n^{3}$
@@
There is one significant weakness of Algorithm 4. In bfloat16 precision, it can introduce numerical errors. Our intuition for why this happens is as follows. Let $\bm{X}=\bm{U}\bm{\Sigma}\bm{V}^{\top}$ be the SVD. For large $T$, $(p_{T}\circ\cdots p_{1})(\bm{X})=\bm{X}\bm{Q}_{T}\approx\operatorname*{polar}(\bm{X})=\bm{U}\bm{V}^{\top}$. Thus, $\bm{Q}_{T}\approx\bm{V}^{\top}\bm{\Sigma}^{-1}\bm{V}$. When $\bm{X}$ has very small singular values and the floating point precision is very low, instantiating $\bm{Q}_{T}$ may be unstable. To mitigate this issue, we use a restarting strategy. Notice that the issue arises only for large $T$, for which $(p_{T}\circ\cdots\circ p_{1})(\epsilon)\approx 1$. Limiting ourselves to $T=3$ iterations improves the conditioning of $\bm{Q}_{T}$ because $(p_{T}\circ\cdots\circ p_{1})(\epsilon)\ll 1$ Thus, to compute $T=6$ iterations, we can apply Algorithm 4 first to the first three polynomials, then again to the last three polynomials. Note that restarting Algorithm 4 after every iteration is exactly the same as the baseline method. This approach provides a tunable hyperparameter—the number of iterations we apply before restarting Algorithm 4—allowing us to find a balance between the fully stable but slow baseline on the one hand, and a very fast but numerically risky method on the other hand.
@@
Figure 8 shows that using Algorithm 4 can dramatically improve runtime on the GPU when the aspect ratio is large enough. As expected, using Algorithm 4 for many iterations significantly reduces the dependence of the runtime on the aspect ratio. Running six iterations of a degree-5 polynomial method when $\alpha=4$ (as with the linear transformations in each MLP block of a transformer) we obtain almost a 2x speedup, and when $\alpha=32$, we obtain nearly a 10x speedup. If we restart every three iterations, the trend is the same but the runtime savings are somewhat smaller.
@@
Figure 8: Effects of using Algorithm 4 on runtime on a GPU. We run $T=6$ iterations of a degree-5 polynomial method on matrices with various dimensions $n$ and aspect ratios $\alpha$. Restart interval $=6$ is Algorithm 4, restart interval $=1$ is equivalent to the baseline (that is, not using Algorithm 4), and restart interval $=3$ is an intermediate method that calls Algorithm 4 once to do the first three iterations and again to do the last three iterations for greater stability. When $\alpha\gg 1$, increasing the restart interval significantly reduces the runtime.
Preliminary experiments using Algorithm 4 with Muon in bfloat16 were not successful. The training and validation losses were sometimes significantly higher than before, even accounting for the runtime savings of Algorithm 4. While this technique may already be applicable in float32 or for architectures whose weight matrices have larger aspect rations, more work is needed to make it practical for general deep learning applications.
@@
If these problems can be mitigated, the speed afforded by Algorithm 4 suggests a potentially beneficial change in the way Muon is applied to transformers. Each multi-head attention layer contains four square weight matrices $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}$ and $\bm{W}_{O}\in\mathbb{R}^{d\times d}$. The orthogonalization step of Muon is either applied separately to these four matrices or else to $[\bm{W}_{Q}\mid\bm{W}_{K}\mid\bm{W}_{V}]$ and $\bm{W}_{O}$, since typical implementations of multi-head attention store the weights in this concatenated form. However, we believe it is natural to consider each of these four weight matrices to be a concatenation of many smaller linear transformations, each corresponding to a single attention head. If $H$ is the number of heads, each of these smaller matrices has size $d\times\frac{d}{H}$; that is, they have aspect ratio $\alpha=H$. Since typical transformers like GPT-3 can have as many as $96$ heads, this version of Muon can yield huge savings in the runtime of each step. We leave it to future work to examine whether this version of Muon also enjoys fast convergence.
@@ Appendix G Code for Constructing Polynomials of Polar Express
The following code gives a Python implementation of the offline stage of Algorithm 2. This code was used to construct the coefficients of the polynomials given in (LABEL:eq:coeffs), which in turn were used in our Muon experiments (Section 5.2). It uses $\ell=10^{-3}$ and $u=1$ by default. It incorporates Algorithm 3 and the numerical tweaks described in Section 4.4

### v3

@@ Appendix D Remez algorithm
  $\begin{bmatrix}a_{k}\\ b_{k}\\ c_{k}\\ E_{k}\end{bmatrix}=\begin{bmatrix}\ell&\ell^{3}&\ell^{5}&1\\ q_{k}&q_{k}^{3}&q_{k}^{5}&-1\\ r_{k}&r_{k}^{3}&r_{1}^{5}&1\\ u&u^{3}&u^{5}&-1\end{bmatrix}^{-1}\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$
  $q_{k+1}=\sqrt{\frac{-3b_{k}-\sqrt{9b_{k}^{2}-20a_{k}c_{k}}}{10c_{k}}},\quad r_{k+1}=\sqrt{\frac{-3b_{k}+\sqrt{9b_{k}^{2}-20a_{k}c_{k}}}{10c_{k}}}$
@@ Appendix E Initialization for Matrices with Large Spectral Gaps
In Section 4, we constructed a sequence of polynomials that is adapted to the range of the singular values $\ell,u]$. Assuming nothing else about the input, these polynomials are optimal since they provide a good approximation to $1$ across the entire interval. However, in many applications, the spectrum has large gaps; that is, there are several large outlying singular values that are well-separated from the rest. For these matrices, it is not necessary for the polynomial to be accurate on the entire interval $\ell,u]$, only on the range of the small singular values plus a few other isolated points. In this section, we take advantage of this structure to accelerate our method by preprocessing the matrix to eliminate the largest singular values.
The first step is to find small intervals containing each of these large singular values. To find lower bounds, we use subspace iteration, which is a generalization of the power method that approximates multiple singular values simultaneously. Fix $k$, the number of singular values we wish to eliminate. Letting $\sigma_{1}\geq\cdots\geq\sigma_{n}$ denote the singular values of $\bm{M}$, subspace iteration produces estimates $\tilde{\sigma}_{1}\geq\cdots\geq\tilde{\sigma}_{k}$ satisfying $\sigma_{i}\geq\tilde{\sigma}_{i}$ for all $i\in 1,\ldots,k$. [^7] Let $\bm{Q}_{0}\in\mathbb{R}^{n\times k}$ be a random matrix with orthonormal columns and define $\bm{Q}_{t+1},\bm{R}_{t+1}=\mathtt{qr}\left(\bm{M}^{\top}\bm{M}\bm{Q}_{t}\right)$, where $\mathtt{qr}$ is the QR decomposition. Subspace iteration outputs the singular values $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{k}$ of $\bm{M}\bm{Q}_{T}$, $\tilde{\sigma}_{1},\ldots,\tilde{\sigma}_{k}$. By the Cauchy interlacing theorem, $\tilde{\sigma}_{k}\leq\sigma_{k}$. To find upper bounds on each $\sigma_{i}$, we can use the fact that $\|\bm{M}\|_{\text{F}}^{2}=\sum_{j=1}^{n}\sigma_{j}^{2}$ as follows:
@@
The second step is to find an odd polynomial that well-approximates the constant function on each of these intervals and on the tail simultaneously. For simplicity, we treat only the $k=1$ case here. Assume that $\bm{M}$ is normalized to $\|\bm{M}\|_{\text{F}}=1$ and let $z=\tilde{\sigma}_{1}$ be the lower bound produced by subspace iteration (which reduces to the power method in this case). Then (20) gives $\sigma_{1}\in[z,1]$ and $\sigma_{2},\ldots,\sigma_{n}\leq\sqrt{1-z^{2}}$. Assume that these intervals do not overlap, that is, $\sqrt{1-z^{2}}\leq z\Leftrightarrow z\geq 1/\sqrt{2}$. Then we construct the unique odd cubic polynomial $p(x)=ax+bx^{3}$ that satisfies $p(\sqrt{1-z^{2}})=1$ and $p(z)=1$ by setting
@@
Because $p(0)=0$ and $p$ has at most one local extremum on $\mathbb{R}_{\geq 0}$, these conditions immediately guarantee that $p$ is concave-increasing on $0,\sqrt{1-z^{2}}]$, so it must lie above the line $x\mapsto x/\sqrt{1-z^{2}}$. Furthermore, $p$ is decreasing on $\sigma_{1},1]$, so it maps $\sigma_{1}\in[z,1]$ to $[p(1),1]$. By minimizing $p(1)$ over all valid $z$ (that is, over the interval $z\in[1/\sqrt{2},1]$), one can further show that $p(1)>1/\sqrt{2}$, so $\sigma_{1}$ cannot be decreased very much by applying $p$. Thus, the largest singular value of $p(\bm{M})$ is still at most $1$, while the smaller singular values have increased by a potentially large factor of $1/\sqrt{1-z^{2}}$. When there is a large outlying singular value, $z$ is close to $1$ and this initialization scheme makes much more progress than a standard iteration of PolarExpress would have.
@@ Appendix F Fast Polynomial Iteration for Rectangular Matrices
  $\bm{R}_{t}=\bm{Q}_{t-1}^{\top}\bm{Y}\bm{Q}_{t-1}$ $\triangleright$ $2n^{3}$
  $\bm{Q}_{t}=\bm{Q}_{t-1}h_{t}(\bm{R}_{t})$ $\triangleright$ Horner’s rule: $\deg(h_{t})\cdot n^{3}$
@@
Algorithm 4 can introduce numerical errors, especially when working in a low precision format like bfloat16. We identify two sources of numerical trouble and propose remedies for each. The first is due to the ill-conditioning of $\bm{X}$. Let $\bm{X}=\bm{U}\bm{\Sigma}\bm{V}^{\top}$ be the SVD. For large $T$, $(p_{T}\circ\cdots p_{1})(\bm{X})=\bm{X}\bm{Q}_{T}\approx\operatorname*{polar}(\bm{X})=\bm{U}\bm{V}^{\top}$. Thus, $\bm{Q}_{T}\approx\bm{V}^{\top}\bm{\Sigma}^{-1}\bm{V}$. When $\bm{X}$ has very small singular values and the floating point precision is very low, instantiating $\bm{Q}_{T}$ may be unstable. To mitigate this issue, we use a restarting strategy. Notice that the issue arises only for large $T$, for which $(p_{T}\circ\cdots\circ p_{1})(\epsilon)\approx 1$. Limiting ourselves to $T=3$ iterations improves the conditioning of $\bm{Q}_{T}$ because $(p_{T}\circ\cdots\circ p_{1})(\epsilon)\ll 1$. Thus, to compute $T>3$ iterations, we begin with $\bm{X}_{0}$ and apply Algorithm 4 with the first three polynomials, producing $\bm{X}_{3}$. When then apply Algorithm 4 again with the next three polynomials to $\bm{X}_{3}$, producing $\bm{X}_{6}$, and so on. As $\bm{X}_{t}$ approaches convergence, its conditioning improves and we may no longer need to restart at all. Note that restarting Algorithm 4 after every iteration is exactly the same as the baseline method.
Second, while the matrix $\bm{Y}$ is positive definite in exact arithmetic, numerical round-off can introduce spurious negative eigenvalues that cause the method to diverge to infinity. To combat this issue, we instead set $\bm{Y}=\bm{X}^{\top}\bm{X}+10^{-3}\bm{I}$ during the first application of Algorithm 4. (We also normalize by $\|\bm{X}\|_{\text{F}}+10^{-3}$ instead of $\|\bm{X}\|_{\text{F}}$.) In subsequent restarts of Algorithm 4, we set $\bm{Y}=\bm{X}^{\top}\bm{X}$ as before. This is akin to slightly increasing each of the singular values of $\bm{X}$, but it does not change the polar factor of $\bm{X}$. Thus, while the output will be slightly different in the early iterations, the algorithm still converges to the correct answer.
@@
Figure 9 shows that using Algorithm 4 can significantly improve runtime on the GPU when the aspect ratio is large enough. As expected, using Algorithm 4 for many iterations significantly reduces the dependence of the runtime on the aspect ratio. Running six iterations of a degree-5 polynomial method when $\alpha=4$ (as with the linear transformations in each MLP block of a transformer) we obtain almost a 2x speedup, and when $\alpha=32$, we obtain a 5x speedup. If we restart every three iterations, the trend is the same but the runtime savings are somewhat smaller.
@@
Figure 9: Effects of using Algorithm 4 on runtime on a GPU. We run $T=6$ iterations of a degree-5 polynomial method on matrices with various dimensions $n$ and aspect ratios $\alpha$. Restart interval =6=6 is Algorithm 4, restart interval =1=1 is equivalent to the baseline (that is, not using Algorithm 4), and restart interval =3=3 is an intermediate method that calls Algorithm 4 once to do the first three iterations and again to do the last three iterations for greater stability. When $\alpha\gg 1$, increasing the restart interval significantly reduces the runtime.
F.1  Application to Muon
@@
If these problems can be mitigated, the speed afforded by Algorithm 4 suggests an improvement in the way Muon is applied to transformers. In sum, the idea is to replace one large matrix with a small aspect ratio by many smaller matrices with large aspect ratios and apply Algorithm 4 to all of them in parallel. Each multi-head attention layer contains four square weight matrices $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}$ and $\bm{W}_{O}\in\mathbb{R}^{d\times d}$. The orthogonalization step of Muon is either applied separately to these four matrices or else to $\bm{W}_{Q}\mid\bm{W}_{K}\mid\bm{W}_{V}]$ and $\bm{W}_{O}$, since typical implementations of multi-head attention store the weights in this concatenated form. However, we believe it is natural to consider each of these four weight matrices to be a concatenation of many smaller linear transformations, each corresponding to a single attention head. If $H$ is the number of heads, each of these smaller matrices has size $d\times\frac{d}{H}$; that is, they have aspect ratio $\alpha=H$. The gradient matrices of $\bm{W}_{Q}\mid\bm{W}_{K}\mid\bm{W}_{V}]$ and $\bm{W}_{O}$ can be reshaped into 3-tensors in which each slice is one of these smaller matrices. Since typical transformers like GPT-3 can have as many as $96$ heads, this variation of Muon has the potential to reduce the runtime.
We use this idea to train a GPT-Small model on FineWeb1B. We compare four conditions:
  1. 1.
    The baseline approach used in the rest of this paper
  2. 2.
    Splitting up the gradient matrices of $\bm{W}_{Q}\mid\bm{W}_{K}\mid\bm{W}_{V}]$ and $\bm{W}_{O}$ by head and applying Muon to each piece, as described above
  3. 3.
    Using Algorithm 4, restarted after three iterations
  4. 4.
    Splitting by head and using Algorithm 4
We used Polar Express with weight decay of $0.1$ for all conditions and swept learning rates $0.003,0.005,0.01$. Otherwise, all hyperparameters were the same as in Section 5.2.
Our results showed that these changes had a negligible effect in this setting. They did not affect the optimization quality. Compared to the baseline, splitting by heads actually reduced the final loss slightly from 3.59 to 3.55; using Algorithm 4 increased the loss very slightly, from 3.59 to 3.60 when not splitting by head, and from 3.55 to 3.56 when we did split. However, the runtimes of all 12 runs were nearly identical, showing that at this scale, the FLOP savings of Algorithm 4 is not beneficial. The embedding size of GPT-Small is just $768$. These techniques may be more impactful when using a larger model. It may also have more impact outside of deep learning, where Polar Express would be run for more than the $5$ iterations used in our experiments. We leave exploration of these settings to future work.
@@ Appendix G Code for Constructing Polynomials of Polar Express
The following code gives a Python implementation of the offline stage of Algorithm 2. This code was used to construct the coefficients of the polynomials given in (1), which in turn were used in our Muon experiments (Section 5.2). It uses $\ell=10^{-3}$ and $u=1$ by default. It incorporates Algorithm 3 and the finite precision considerations described in Section 4.4.